# STF: Shallow-Level Temporal Feedback to Enhance Spiking Transformers

### v1

@@ Abstract
Transformer-based Spiking Neural Networks (SNNs) suffer from a great performance gap compared to floating-point Artificial Neural Networks (ANNs) due to the binary nature of spike trains. Recent efforts have introduced deep-level feedback loops to transmit high-level semantic information to narrow this gap. However, these designs often span multiple deep layers, resulting in costly feature transformations, higher parameter overhead, increased energy consumption, and longer inference latency. To address this issue, we propose Shallow-level Temporal Feedback (STF), a lightweight plug-and-play module for the encoding layer, which consists of Temporal-Spatial Position Embedding (TSPE) and Temporal Feedback (TF). Extensive experiments show that STF consistently improves performance across various Transformer-based SNN backbones on static datasets, including CIFAR-10, CIFAR-100, and ImageNet-1K, under different spike timestep settings. Further analysis reveals that STF enhances the diversity of the spike patterns, which is key to performance gain. Moreover, evaluations on adversarial robustness and temporal sensitivity confirm that STF outperforms direct coding and its variants, highlighting its potential as a new spike encoding scheme for static scenarios. Our code will be released upon acceptance.
@@ 1 Introduction
Building on this insight, we propose a Shallow-level Temporal Feedback (STF) module that can be seamlessly integrated into the encoding layer of existing models. As illustrated in Figure 1(b), STF comprises Temporal-Spatial Position Embedding (TSPE) and Temporal Feedback (TF), which together introduce temporal feedback efficiently without compromising the spike-driven nature of SNNs. We evaluate the impact of STF on several mainstream Transformer-based SNN architectures, including Spikformer (Zhou et al. 2023), Spike-driven Transformer (SDT) (Yao et al. 2023), and QKFormer (Zhou et al. 2024). Experimental results demonstrate that STF consistently improves accuracy across different spike timesteps on static datasets such as CIFAR-10, CIFAR-100, and ImageNet-1K. Notably, on the large-scale ImageNet-1K dataset, STF improves accuracy by 0.75%, with only a 0.32% increase in parameters, 7.2% more energy consumption, and 6.33% higher inference latency. Further analysis, including visualizations (Figure 2), quantitative comparisons based on spike entropy (Table 1), and theoretical analysis (Section 4.3), reveals that the key factor behind STF’s performance gains lies in its ability to enhance the diversity of the spike patterns. This enables richer spike-based representations, which are critical for static visual tasks. In addition, we conduct a detailed evaluation of STF against direct coding and its variants, focusing on model performance, robustness, and temporal information preservation. The comparative experiments suggest that STF represents a promising new spike encoding scheme, with the potential to replace the long-standing dominance of direct coding in static data scenarios. Our main contributions can be summarized as follows:
  • We propose the STF module, a plug-and-play temporal feedback mechanism applied at the encoding layer. It preserves the spike-driven nature of SNNs by introducing feedback at low-dimensional feature layers and directly modulating spiking neurons. STF consistently improves model performance under varying spike timesteps across multiple datasets, including CIFAR-10, CIFAR-100, and ImageNet-1K, while incurring only modest increases in model parameters, energy consumption, and inference latency.
  • Through detailed visualizations, quantitative comparisons, and theoretical analysis, we identify that the key to STF’s performance improvement lies in its ability to enhance the diversity of the spike patterns, thereby improving spike-based representation. Besides, we also systematically evaluate four STF variants and identify the optimal configuration.
@@ 2.1 Transformer-based SNNs
Due to their strong performance across a wide range of tasks, Transformer-based SNNs have garnered increasing attention from researchers, leading to continuous architectural innovations. Spikformer (Zhou et al. 2023) was the first to integrate the Transformer architecture with SNNs, replacing the softmax operation with binary spike-based representations to effectively reduce energy consumption. Spike-driven Transformer (Yao et al. 2023, 2024) further advanced the paradigm by substituting the dot-product attention with Hadamard product operations and redesigning residual connections based on membrane potentials to establish a fully spike-driven framework. QKFormer (Zhou et al. 2024) and SpikeRWKV (Zhu et al. 2025a) introduced linear attention mechanisms to further reduce computational complexity, while E-Spikeformer (Yao et al. 2025) adopted integer-based training approaches to enhance both training efficiency and model performance. Despite these advances, most existing Transformer-based SNNs still rely on ANN-inspired designs, neglecting bio-inspired principles in architectural design that may unlock further performance gains.
@@ 2.2 Feedback Mechanisms in SNNs
Originally rooted in cybernetics (Wiener 2019) and later adopted in neuroscience (Lamme and Roelfsema 2000; Friston 2005), feedback mechanisms provide strong self-adaptive regulation and have been widely integrated into SNNs. BackEISNN (Zhao, Zeng, and Li 2022) and RSNN (Xu et al. 2024) applied recurrent feedback to all neurons and introduced additional hidden states. BRP (Zhang et al. 2021) used cortical reward signal transmission to feed label information back to all hidden layers, enhancing accuracy through local pseudo-gradients. STSF (He et al. 2025) and STNet (Duan et al. 2025) adopted spatio-temporal feedback across cortical hierarchies to mitigate feature degradation and temporal loss. Besides, SpikingVTG (Bal et al. 2024) was the first to utilize high-layer firing rates to modulate attention in intermediate layers, improving performance in video-text temporal grounding. SpiLiFormer (Zheng et al. 2025) and TDFormer (Zhu et al. 2025b) incorporated lateral inhibition and top-down modulation, respectively, to address attention dispersion. However, these approaches primarily rely on structurally deep-level feedback loops, where high-level semantic information is propagated back to multiple or even all hidden layers through high-dimensional feature transformations. In contrast, the shallow-level and temporal-wise feedback mechanism has not yet been thoroughly investigated.
@@ 3.1 Spiking Neuron
where $H[t]$ denotes the intermediate membrane state at time $t$, which is influenced by the input current $I[t]$, the previous membrane potential $U[t-1]$, and the reset voltage $U_{\text{reset}}$. The output spike $S[t]$ is generated when $H[t]$ exceeds the threshold $U_{\text{th}}$, determined by the Heaviside function $\Theta(\cdot)$, after which the membrane potential is reset to $U_{\text{reset}}$.
@@
Table 1: Performance comparison of our method on 333 backbones: Spikformer, Spike-driven Transformer (SDT), and QKFormer, on the CIFAR datasets across various time steps. Bold indicates the best result, and underlined denotes the second-best.
@@ 3.2 Shallow-Level Temporal Feedback Module
To effectively capture temporal and spatial dynamics in spike trains, position embedding plays a crucial role in SNNs (Lv et al. 2024). We represent the input image sequence as $I\in\mathbb{R}^{T\times C\times H\times W}$, where $T$ is the number of spike timesteps, and $C$, $H$, and $W$ refer to the number of channels, image height, and image width, respectively. In STF-1 and STF-3 configurations, as shown in Figure 1(c), the TSPE mechanism is formulated as follows:
@@
where $X_{\text{TPE}}\in\mathbb{R}^{T\times C\times H\times W}$ represents a spatio-temporal positional embedding, and $W_{\text{ConvBN}}$ denotes the weights of a 2D convolution layer followed by batch normalization. Motivated by the natural similarity between spike trains and video signal processing, a 3D trigonometric positional strategy is adopted to initialize $X_{\text{TPE}}$, following common practice in video transformers (Liu et al. 2022; Liang et al. 2024). To explore the optimal location for incorporating TSPE, two additional variants, STF-2 and STF-4, are introduced. In these cases, the embedding formulation is modified as follows:
@@
To maintain the spike-driven nature of the network, TF directly feeds the spike output from the previous time step $S[t-1]$ into the input at the current time step $S[t]$. This forms a feedback loop that captures temporal dependencies. In contrast to conventional recurrent SNNs (Zhao, Zeng, and Li 2022; Xu et al. 2024), TF avoids the use of auxiliary hidden states or non-spiking activations. The feedback operation in STF-1 and STF-2, as illustrated in Figure 1(c), is formulated as:
@@
where all variables follow the definitions provided in Equations 13.
@@ 4 Experiments
E-SpikeFormer E-SpikeFormer-12-768 224 173.0 35.6   4$\star$ 84.70
E-SpikeFormer-12-768 224 173.0 54.7   8$\star$ 85.10
   E-SpikeFormer-12-768** 384 173.0 -   8$\star$ 86.20
SDT SDT-6-512 224 23.37 3.56 4  72.41$\dagger$
@@
Table 2: Performance comparison on ImageNet-1K. Numbers marked with $\dagger$ indicate our implementation based on publicly available code. * and **∗ ∗ indicate inference resolutions of $288^{2}$ and $384^{2}$, respectively. $\star$ represents a quantized training method (Yao et al. 2025) using integer-only activations, which are later mapped to spike trains for inference.
@@ 4.2 Effectiveness of STF on Image Classification Tasks
To answer RQ1, we train various models from scratch and evaluate the performance improvements introduced by the STF module across three datasets and three Transformer-based SNN backbones. As shown in Table 1, STF consistently outperforms the corresponding baselines on both CIFAR-10 and CIFAR-100 across four different timesteps. Specifically, on CIFAR-100, STF improves performance by 1.77% with Spikformer at $T=2$ and by 1.04% with SDT at $T=4$.
We further evaluate STF on the more challenging ImageNet-1K dataset to assess its performance improvement over SDT and QKFormer, along with the additional parameter and energy consumption overhead. As shown in Table 2, STF improves accuracy by 0.6% to 0.75%. Specifically, for SDT with the timestep of $T=4$, STF achieves a 0.6% accuracy gain while introducing only a 0.17% increase in parameters and a 12.08% increase in energy consumption. Similarly, for QKFormer with $T=4$ and a standard input resolution of 224224224, STF improves accuracy by 0.75% with just a 0.32% increase in parameters and approximately 7.2% more energy consumption. Furthermore, STF allows QKFormer, which uses only 37.5% of the parameters of the best-performing E-Spikformer, to reach comparable accuracy (86.2%) with only half the number of timesteps.
@@ 4.3 Analyzing the Performance Gains of STF
To answer RQ2, we analyze the distribution of the spike train of the same LIF neuron in the encoding layer, with and without STF, in the CIFAR datasets. Specifically, we visualize the output patterns under STF and the direct coding scheme for comparison. As shown in Figure 2, taking $T=4$ as an example, STF alleviates the imbalance in the distribution of the spike patterns to varying degrees between different main SNN models. It significantly increases the number of activated spike patterns and successfully covers all 16 possible binary combinations. A more diverse spike pattern indicates greater information capacity, facilitating more effective information propagation through the network and ultimately enhancing model representation and performance. Therefore, we consider that the improved diversity of the spike patterns is a key contributor to STF performance gains.
To quantitatively measure this property, we adopt spike entropy as a diversity metric, inspired by previous work in neuroscience and information theory (Luczak 2024). The formal definition is provided in Appendix H. As reported in Table 1, STF consistently improves the diversity of the spike patterns across all models and datasets, with greater gains at higher spike timesteps.
We further provide a theoretical analysis to explain why STF can improve the diversity of the spike patterns. For static datasets, the input image $I$ is duplicated across $T$ spike timesteps to form a sequence $I=[I[1],I[2],...,I[T]]$, where $I[t]=I$ for all $t\in{1,2,...,T}$. The spike generation time $SG_{t}$ of a neuron in the encoding layer without STF can be formulated as follows (the detailed derivation is provided in Appendix G):
@@
This expression indicates that under fixed values of leakage decay factor $\tau$, $U_{\text{th}}$, and a repeated static input $I$, baseline models tend to activate only a limited set of specific spike patterns. In contrast, STF incorporates both the current input $I[t]$ and the previous spike output $S[t{-}1]$ into Equation 7, allowing $SG_{t}$ to be triggered over a broader temporal range, rather than being constrained by the constant input $I$. This effectively expands the expressive space of spike patterns, leading to enhanced diversity in spike patterns. In addition, we conduct detailed ablation studies on TSPE and TF, as presented in Appendix J. The results demonstrate that both components significantly contribute to improving the diversity of the spike patterns and overall model performance, with TF exhibiting a more pronounced effect.
@@ 4.4 Comparison with Existing Spiking Encoding Schemes
To answer RQ3, we evaluate two representative spike encoding variants, GAC (Qiu et al. 2024) and IMP (Shen et al. 2024), in terms of model performance and the diversity of the spike patterns on the CIFAR datasets. As shown in the spike pattern visualizations in Figure 2 and the results in Table 1, GAC brings performance improvements in some models. However, it exhibits a strong preference for specific spike patterns, similar to the direct coding scheme, resulting in a concentrated activation distribution. In contrast, IMP demonstrates a more pronounced effect in increasing the diversity of the spike patterns, which contributes to performance gains. Nevertheless, it still fails to activate the full range of possible spike combinations, indicating limited diversity. STF, on the other hand, consistently outperforms both GAC and IMP in terms of the diversity of the spike patterns, leading to further improvements in model accuracy. Although STF yields slightly lower spike entropy values than IMP under certain timestep configurations, its advantage becomes increasingly evident as the number of timesteps grows, demonstrating better scalability and stability.

### v2

@@ Abstract
Transformer-based Spiking Neural Networks (SNNs) suffer from a great performance gap compared to floating-point Artificial Neural Networks (ANNs) due to the binary nature of spike trains. Recent efforts have introduced deep-level feedback loops to transmit high-level semantic information to narrow this gap. However, these designs often span multiple deep layers, resulting in costly feature transformations, higher parameter overhead, increased energy consumption, and longer inference latency. To address this issue, we propose Shallow-level Temporal Feedback (STF), a lightweight plug-and-play module for the encoding layer, which consists of Temporal-Spatial Position Embedding (TSPE) and Temporal Feedback (TF). Extensive experiments show that STF consistently improves performance across various Transformer-based SNN backbones on static datasets, including CIFAR-10, CIFAR-100, and ImageNet-1K, under different spike timestep settings. Further analysis reveals that STF enhances the diversity of spike patterns, which is key to performance gain. Moreover, evaluations on adversarial robustness and temporal sensitivity confirm that STF outperforms direct coding and its variants, highlighting its potential as a new spike encoding scheme for static scenarios. Our code will be released upon acceptance.
@@ 1 Introduction
Building on this insight, we propose a Shallow-level Temporal Feedback (STF) module that can be seamlessly integrated into the encoding layer of existing models. As illustrated in Figure 1(b), STF comprises Temporal-Spatial Position Embedding (TSPE) and Temporal Feedback (TF), which together introduce temporal feedback efficiently without compromising the spike-driven nature of SNNs. We evaluate the impact of STF on several mainstream Transformer-based SNN architectures, including Spikformer (Zhou et al. 2023), Spike-driven Transformer (SDT) (Yao et al. 2023), and QKFormer (Zhou et al. 2024). Experimental results demonstrate that STF consistently improves accuracy across different spike timesteps on static datasets such as CIFAR-10, CIFAR-100, and ImageNet-1K. Notably, on the large-scale ImageNet-1K dataset, STF improves accuracy by 0.75%, with only a 0.32% increase in parameters, 7.2% more energy consumption, and 6.33% higher inference latency. Further analysis, including visualizations (Figure 2), quantitative comparisons based on spike entropy (Table 1), and theoretical analysis (Section 4.3), reveals that the key factor behind STF’s performance gains lies in its ability to enhance the diversity of spike patterns. This enables richer spike-based representations, which are critical for static visual tasks. In addition, we conduct a detailed evaluation of STF against direct coding and its variants, focusing on model performance, robustness, and temporal information preservation. The comparative experiments suggest that STF represents a promising new spike encoding scheme, with the potential to replace the long-standing dominance of direct coding in static data scenarios. Our main contributions can be summarized as follows:
  • We propose the STF module, a plug-and-play temporal feedback mechanism applied at the encoding layer. It preserves the spike-driven nature of SNNs by introducing feedback at low-dimensional feature layers and directly modulating spiking neurons. STF consistently improves model performance under varying spike timesteps across CIFAR-10, CIFAR-100, and ImageNet-1K, while incurring only modest increases in model parameters, energy consumption, and inference latency.
  • Through detailed visualizations, quantitative comparisons, and theoretical analysis, we identify that the key to STF’s performance improvement lies in its ability to enhance the diversity of spike patterns, thereby improving spike-based representation. Besides, we systematically evaluate four architectural variants of STF and determine the optimal configuration.
@@ 2.1 Transformer-based SNNs
Due to their strong performance across a wide range of tasks, Transformer-based SNNs have garnered increasing attention from researchers, leading to continuous architectural innovations. Spikformer (Zhou et al. 2023) was the first to integrate the Transformer architecture with SNNs, replacing the softmax operation with binary spike-based representations to effectively reduce energy consumption. Spike-driven Transformer (Yao et al. 2023, 2024) further advanced the paradigm by substituting the dot-product attention with Hadamard product operations and redesigning residual connections based on membrane potentials to establish a fully spike-driven framework. QKFormer (Zhou et al. 2024) and SpikeGPT (Zhu et al. 2025a) introduced linear attention mechanisms to further reduce computational complexity, while E-SpikeFormer (Yao et al. 2025) adopted integer-based training approaches to enhance both training efficiency and model performance. Despite these advances, most existing Transformer-based SNNs still rely on ANN-inspired designs, neglecting bio-inspired principles in architectural design that may unlock further performance gains.
@@ 2.2 Feedback Mechanisms in SNNs
Originally rooted in cybernetics (Wiener 2019) and later adopted in neuroscience (Lamme and Roelfsema 2000), feedback mechanisms provide strong self-adaptive regulation and have been widely integrated into SNNs. BackEISNN (Zhao, Zeng, and Li 2022) and RSNN (Xu et al. 2024) applied recurrent feedback to all neurons and introduced additional hidden states. BRP (Zhang et al. 2021) used cortical reward signal transmission to feed label information back to all hidden layers, enhancing accuracy through local pseudo-gradients. STSF (He et al. 2025) and STNet (Duan et al. 2025) adopted spatio-temporal feedback across cortical hierarchies to mitigate feature degradation and temporal loss. Besides, SpikingVTG (Bal et al. 2024) was the first to utilize high-layer firing rates to modulate attention in intermediate layers, improving performance in video-text temporal grounding. SpiLiFormer (Zheng et al. 2025) and TDFormer (Zhu et al. 2025b) incorporated lateral inhibition and top-down modulation, respectively, to address attention dispersion. However, these approaches primarily rely on structurally deep-level feedback loops, where high-level semantic information is propagated back to multiple or even all hidden layers through high-dimensional feature transformations. In contrast, the shallow-level and temporal feedback mechanism has not yet been thoroughly investigated.
@@ 3.1 Spiking Neuron
where $H[t]$ denotes the intermediate membrane state at time $t$, which is influenced by the input current $I[t]$, the previous membrane potential U[t1]U[t-1], and the reset voltage $U_{\text{reset}}$. The output spike $S[t]$ is generated when $H[t]$ exceeds the threshold $U_{\text{th}}$, determined by the Heaviside function $\Theta(\cdot)$, after which the membrane potential is reset to $U_{\text{reset}}$.
@@
Table 1: Performance comparison of our method on 33 backbones: Spikformer, Spike-driven Transformer (SDT), and QKFormer, on the CIFAR datasets across various time steps. Bold indicates the best result, and underlined denotes the second-best.
@@ 3.2 Shallow-Level Temporal Feedback Module
To effectively capture temporal and spatial dynamics in spike trains, position embedding plays a crucial role in SNNs (Lv et al. 2024). In our formulation, we represent the input image sequence as $I\in\mathbb{R}^{T\times C\times H\times W}$, where $T$ is the number of spike timesteps, and $C$, $H$, and $W$ refer to the number of channels, image height, and image width, respectively. In STF-1 and STF-3 configurations, as shown in Figure 1(c), the TSPE mechanism is formulated as follows:
@@
where $X_{\text{TPE}}\in\mathbb{R}^{T\times C\times H\times W}$ represents a spatio-temporal positional embedding, and $W_{\text{ConvBN}}$ denotes the weights of a 2D convolution layer followed by batch normalization. Motivated by the inherent similarity between spike trains and sequential video signal processing, a 3D trigonometric positional strategy is adopted to initialize $X_{\text{TPE}}$, following common practice in video transformers (Liu et al. 2022; Liang et al. 2024). To explore the optimal location for incorporating TSPE, two additional variants, STF-2 and STF-4, are introduced. In these configurations, the embedding formulation is modified as follows:
@@
To maintain the spike-driven nature of the network, TF directly feeds the spike output from the previous time step S[t1]S[t-1] into the input at the current time step $S[t]$. This forms a feedback loop that captures temporal dependencies. In contrast to conventional recurrent SNNs (Zhao, Zeng, and Li 2022; Xu et al. 2024), TF avoids the use of auxiliary hidden states or non-spiking activations. The feedback operation in STF-1 and STF-2, as illustrated in Figure 1(c), is formulated as:
@@
where all variables strictly adhere to the definitions provided in Equations 13.
@@ 4 Experiments
[…]
E-SpikeFormer-12-768 224 173.0 54.7   $\star 8$5.10
   E-SpikeFormer-12-768** 384 173.0 -   $\star 8$6.20
[…]
@@
Table 2: Performance comparison on ImageNet-1K. Numbers marked with $\dagger$ indicate our implementation based on publicly available code. * and ** indicate inference resolutions of $288^{2}$ and $384^{2}$, respectively. $\star$ represents a quantized training method (Yao et al. 2025) using integer-only activations, which are later mapped to spike trains for inference.
@@ 4.2 Effectiveness of STF on Image Classification Tasks
To answer RQ1, we train various models from scratch and evaluate the performance improvements introduced by STF across three datasets and three Transformer-based SNN backbones. As shown in Table 1, STF consistently outperforms the corresponding baselines on both CIFAR-10 and CIFAR-100 across four different timesteps. Specifically, on CIFAR-100, STF improves performance by 1.77% with Spikformer at $T=2$ and by 1.04% with SDT at $T=4$.
We further evaluate STF on the more challenging ImageNet-1K dataset to assess its performance improvement over SDT and QKFormer, along with the additional parameter and energy consumption overhead. As shown in Table 2, STF improves accuracy by 0.6% to 0.75%. Specifically, for SDT with the timestep of $T=4$, STF achieves a 0.6% accuracy gain while introducing only a 0.17% increase in parameters and a 12.08% increase in energy consumption. Similarly, for QKFormer with $T=4$ and a standard input resolution of $224$, STF improves accuracy by 0.75% with just a 0.32% increase in parameters and approximately 7.2% more energy consumption. Furthermore, STF allows QKFormer, which uses only 37.5% of the parameters of the best-performing E-SpikeFormer, to reach comparable accuracy (86.2%) with only half the number of timesteps.
@@ 4.3 Analyzing the Performance Gains of STF
To answer RQ2, we analyze the distribution of spike trains from the same LIF neuron in the encoding layer, both with and without STF, on the CIFAR datasets. Specifically, we visualize the output patterns under STF and the direct coding scheme for comparison. As shown in Figure 2, taking $T=4$ as an example, STF alleviates the imbalance in the distribution of spike patterns to varying degrees between different main SNN models. It significantly increases the number of activated spike patterns and successfully covers all 16 possible binary combinations. A more diverse spike pattern indicates greater information capacity, facilitating more effective information propagation through the network and ultimately enhancing model representation and performance. Therefore, we consider that the improved diversity of spike patterns is a key contributor to STF performance gains.
To quantitatively measure this property, we adopt spike entropy as a diversity metric, inspired by previous work in neuroscience and information theory (Luczak 2024). The formal definition is provided in Appendix H. As reported in Table 1, STF consistently improves the diversity of spike patterns across all models and datasets, with greater gains at higher spike timesteps.
We further provide a theoretical analysis to explain why STF can improve the diversity of spike patterns. For static datasets, the input image $I$ is duplicated across $T$ spike timesteps to form a sequence I=[I[1],I[2],,I[T]]I=[I[1],I[2],...,I[T]], where $I[t]=I$ for all $t\in{1,2,...,T}$. The spike generation time $SG_{t}$ of a neuron in the encoding layer without STF can be formulated as follows (the detailed derivation is provided in Appendix G):
@@
This expression indicates that under fixed values of leakage decay factor $\tau$, $U_{\text{th}}$, and a repeated static input $I$, baseline models tend to activate only a limited set of specific spike patterns. In contrast, STF incorporates both the current input $I[t]$ and the previous spike output S[t1]S[t{-}1] into Equation 7, allowing $SG_{t}$ to be triggered over a broader temporal range, rather than being constrained by the constant input $I$. This effectively expands the expressive space of spike patterns, leading to enhanced diversity in spike patterns. In addition, we conduct detailed ablation studies on TSPE and TF, as presented in Appendix J. The results demonstrate that both components significantly contribute to improving the diversity of spike patterns and overall model performance, with TF exhibiting a more pronounced effect.
@@ 4.4 Comparison with Existing Spiking Encoding Schemes
To answer RQ3, we evaluate two representative spike encoding variants, GAC coding (Qiu et al. 2024) and IMP coding (Shen et al. 2024), in terms of model performance and the diversity of spike patterns on the CIFAR datasets. As illustrated by the spike pattern visualizations in Figure 2 and the results in Table 1, GAC provides performance improvements for certain models. However, it exhibits a strong preference for specific spike patterns, similar to the direct coding scheme, resulting in a concentrated activation distribution. In contrast, IMP demonstrates a more pronounced effect in increasing the diversity of spike patterns, which contributes to performance gains. Nevertheless, it still fails to activate the full range of possible spike combinations, indicating limited diversity. STF, on the other hand, consistently outperforms both GAC and IMP in terms of the diversity of the spike patterns, leading to further improvements in model accuracy. Although STF yields slightly lower spike entropy values than IMP under certain timestep configurations, its advantage becomes increasingly evident as the number of timesteps grows, demonstrating better scalability and stability.